asked for this problem but i wrote it incorrectly..
-54/3 = 3/2x+2

Answer by tutorcecilia(2144)

-54/3 = 3/2x+2 [solve for x]
6(-54/3 = (3/2)x+2) [multiply each term by the lowest common denominator=6]
-108=9x+12 [simplify]
-108-12=9x
-120=9x
-120/9=9x/9
-40/3=x
.
checking:
-54/3-(3/2)(-40/3)+2
-18=-18 [checks out]